TrailPacker produces rugged backpacks for outdoor sports (hiking, rock climbing, etc.) Their backpacks are sold at many specialty outdoor stores across the country. The cost of manufacturing and marketing their backpacks, at their normal factory volume of 20,000 backpacks per month, is shown in the table below. TrailPacker sells these backpacks for $50 each. TrailPacker is making a small profit, but they would prefer to increase their Operating Income.

Hint: Fixed costs are shown on a per-unit basis in the table based on normal volume. However, fixed costs as a total do not change when volume changes, so you will need to determine total fixed costs first.

Per Unit

Per Unit

Unit Manufacturing Cost

$ 7.00

Variable Materials

$ 11.00

Variable Labor

$ 6.00

Variable Overhead

$ 10.00

Total Unit Manufacturing Cost

$ 34.00

Unit Marketing Costs

Variable Marketing Costs

$ 2.00

Fixed Marketing Costs

$ 4.00

Total Unit Marketing Costs

$ 6.00

A) TrailPacker wants to understand their basic starting financial data. What is their monthly fixed cost, variable cost per backpack, and contribution margin per backpack? Show your calculations for each.

B) Prepare a one-month Contribution Margin Income Statement for the company using the given financial data at their normal factory volume. Include line items for each type of cost as well as subtotals for the variable and fixed costs.

C) What is the break-even point in units? (Show your calculations.)

D) What is the break-even point in sales dollars? (Show your calculations.)

E) Using a one-month Contribution Margin Income Statement, verify that your calculated break-even volume results in Operating Income of Zero. (Prepare the entire Contribution Margin statement at the break-even level.)
Answer #1

Monthly fixed cost = fixed marketing cost = 4*20,000 =$80,000

Variable cost per backpack = 7+11+6+10+2 =$36 per backpack

Contribution margin per unit = selling price per unit- variable costs per unit

=50-36 =$14 per backpack

B. Contribution margin income statement

Sales revenue 20,000*50 =$1,000,000

Less: variable costs

Manufacturing is 7*20,000 =$140,000

Direct material 11*20,000 =$220,000

Variable labor 6*20,000 =120,000

Variable overhead 10*20,000 =$200,000

Variable marketing cost =2*20,000 =$40,000

Total variable cost =$720,000

Contribution margin =$280,000

Less fixed costs

Fixed marketing cost 80,000

Net operating income =$200,000

C.break even point in units = fixed costs/contribution margin per unit

=80,000/14

= 5714.29 units

D.in sales dollars = fixed costs/contribution margin ratio

Contribution margin ratio = contribution margin/sales

=14/50

= 28%

Break even point = 80,000/28%

=$285,714.29

E. Sales revenue =$285,714.29

Less: variable cost =$205,714.29

Contribution margin =$80,000

Fixed costs =$80,000

Operating income =$0